the mean grade in physics examination was 75 and the standard deviation was 5. Determine the grade of the student whose standard is -0.5



1
Expert's answer
2021-11-16T10:27:59-0500

"\mu=75 \\

\sigma=5"

Standard score z = -0.5

"z = \frac{x- \mu}{\sigma} \\

-0.5 = \frac{x -75}{5} \\

-2.5 = x -75 \\

x = 75-25 = 50"

Answer: 50